OpenStudy (anonymous):

Find the center, the radius, the x intercepts and y intercepts for the circle... x^2 + y^2 -8x +6y-11=0

OpenStudy (anonymous):

x^2 - 8x + y^2 + 6y = 11 x^2 - 8x+ (4)^2 +y^2 + 6y + (3)^2 = 11 = (4)^2 + (3)^2 .....? (x-4)^2 + (y+3)^2 = 28............? Radius= √28...........? Center (4, -3) ..........?